OpenStudy (anonymous):

A research submarine has a 30-cm-diameter window that is 8.6cm thick. The manufacturer says the window can withstand forces up to 1.2×10^6 N . What is the submarine's maximum safe depth in salt water? The pressure inside the submarine is maintained at 1.0 atm.

OpenStudy (anonymous):

Thickness of window is not req'd. Window area = pi.0.3*0.3/4 = 0.07 sq m. Force = 1200000=press..07. Press = 1200000/.07 = 17142857 n/sq m. Press = depth*density , assume seawater density = 1030 kg/cu m. so 1030 *g*depth = 17142857 depth = 17142857/9.81*1030 = 1698 m.

OpenStudy (anonymous):

Thanks, I was stuck on this problem forever. Now I know what I did wrong.